# Espresso: Robust Concept Filtering in Text-to-Image Models

### v2

@@ 3. Problem Statement
Adversary Model. We assume (blackbox) access to a deployed target T2I model (f) to which a client can send an input (p) to generate an image ($x$). Further, f uses some CRT. The goal of the adversary ($\mathcal{A}dv$) is to force f to generate an image containing an unacceptable concept ($x^{u}$) despite the presence of a CRT to suppress it. We give an advantage to $\mathcal{A}dv$ by allowing whitebox access to a local identical copy of f with the CRT for use in designing attacks. This is reasonable as $\epsilon_{\theta}$ and CLIP are publicly available. For filtering-based CRTs, we assume that $\mathcal{A}dv$ has whitebox access to the filter, enabling them to use its loss function in designing the attacks.
@@
Requirements An ideal CRT should be: R1 Effectivein minimizing the generation of images with unacceptable concepts ($x^{u}$); R2 Utility-preserving, maintaining the quality of acceptable images (for fine-tuning CRTs) or not blocking them (for filtering CRTs); and R3 Robustto resist evasion with adversarial prompts ($\texttt{p}^{adv}$).
@@
Limitations of Prior Works. In Section 6, we empirically show the limitations of prior works by evaluating R1, R2, and R3 of different CRTs. We provide a brief summary here. Fine-tuning CRTs (Gandikota et al., 2023a; Kumari et al., 2023a; Zhang et al., 2023b; Heng and Soh, 2023b, b; Gandikota et al., 2024; Kim et al., 2023) modify f thereby explicitly creating a trade-off between effectiveness (R1) and utility (R2). Further, these CRTs do not consider robustness (R3) in their design and hence, are susceptible to evasion by $\mathcal{A}dv$ using $\texttt{p}^{adv}$. Filtering CRTs (Radford et al., 2021; Qu et al., 2023b) detect unacceptable concepts either in p (aka prompt filter) or in $x$ and block them (a.k.a image filter). Since, they do not modify f, they can maintain utility without impacting effectiveness. Prior filtering approaches may not be accurate in detecting unacceptable concepts (impacts R1(Li et al., 2024). They can also be easily evaded (harms R3) ((Rando et al., 2022) and Section 6.2). Further, the current state-of-the-art filter, UD (Qu et al., 2023b), trains specialized classifiers for each concept on substantial data, which limits their generalization to new concepts.
4. Espresso: Robust Filtering CRT
@@
We present Espresso, a robust concept filtering CRT which satisfies all the requirements from Section 3. Espresso uses a classifier to detect images with unacceptable concepts. We identify CLIP as the natural choice for such a classifier as it is (1) pre-trained on a large dataset covering a wide range of concepts, and (2) used across many T2I models, and encodes similar information as seen in these models. Hence, CLIP is a better choice for a filter than training specialized classifiers to identify specific concepts (e.g., (Qu et al., 2023b)).
@@
However, simply using CLIP for Espresso is not sufficient as seen in Stable Diffusion’s filter (Radford et al., 2021) $\texttt{F}_{SD}$. Here, the filter thresholds the cosine similarity between the embeddings of $x$ and each pre-defined unacceptable concept to identify $x^{u}$. Rando et al. (Rando et al., 2022) design adversarial prompts ($\texttt{p}^{adv}$) to evade $\texttt{F}_{SD}$. As $\texttt{F}_{SD}$ uses only the cosine similarity to $c^{u}$, it gives $\mathcal{A}dv$ the freedom to modify the prompt embeddings in any direction while optimizing for $\texttt{p}^{adv}$, to force a misclassification. We address this in Espresso to ensure effectiveness and robustness while maintaining utility by (a) modifying CLIP’s classification objective for R1 and R3, and (b) fine-tuning to further improve Espresso.
@@
Modifying CLIP’s Classification Objective. We modify the objective function of Espresso for filtering by using the cosine similarity to both $c^{a}$ and $c^{u}$ instead of just the cosine similarity to $c^{u}$. Further, as observed in prior fine-tuning CRTs, jointly optimizing for two embeddings yields better effectiveness and utility (Kumari et al., 2023a; Tsai et al., 2024; Gal et al., 2022). Using the modified classification objective, we restrict $\mathcal{A}dv$ to adding noise only along the low-dimensional vector joining $c^{a}$ and $c^{u}$, in the direction of $c^{a}$. This is more challenging for $\mathcal{A}dv$ to evade than $\texttt{F}_{SD}$, as projecting to a lower-dimensional representation can improve robustness (Wang and Liu, 2021; Bhagoji et al., 2018).
@@
(1) $=\text{argmax}_{i}\left(\left\{\frac{\text{exp}{(\text{cos}(\phi_{x}(x),\phi_{p}(c_{i}))/\tau)}}{\sum_{j\in\{a,u\}}\text{exp}(\text{cos}(\phi_{x}(x),\phi_{p}(c_{j}))/\tau)}\right\}_{i\in\{a,u\}}\right)$
where $\tau=\frac{1}{100}$ is the default temperature parameter used in CLIP. Further, the cosine similarity is equivalent to the scaled Euclidean distance between the embeddings (Liang et al., 2022):
@@
$\implies\text{cos}(\texttt{$x$},c)=\frac{-1}{2}(\overline{\phi_{x}(\texttt{$x$})}-\overline{\phi_{p}(c)})^{T}(\overline{\phi_{x}(\texttt{$x$})}-\overline{\phi_{p}(c)})+1$
Hence, by using CLIP as a filter, we project the image embeddings onto the vector between $c^{u}$ and $c^{a}$ and classify it based on its distance to both.
Fine-tuning. We conjecture that we can further improve Espresso by maximizing the distance between $\phi_{p}(\texttt{p}^{u})$ and $\phi_{p}(\texttt{p}^{a})$. In other words, we can fine-tune F to minimize the following loss function:
(2) $L_{\textsc{Espresso}}=-||\overline{\phi_{p}(\texttt{p}^{u})-\phi_{p}(\texttt{p}^{a})}||_{2}$
For the case where $\texttt{p}^{u}$ and $\texttt{p}^{a}$ are opposites (e.g., they contain violence and peaceful), the above objective function might result in a low correlation between $\phi_{p}(\texttt{p})$ and the corresponding $\phi_{x}(\texttt{$x$})$. This might result in poor utility. We account for this by using the following objective function instead:
(3) $\begin{split}\mathcal{L}_{\textsc{Espresso}}=\alpha_{\text{aa}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{aa})-\alpha_{\text{ua}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{ua})\\ +\alpha_{\text{uu}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{uu})-\alpha_{\text{au}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{au})\\ +\alpha_{\text{uu-t}}\mathcal{L}_{MSE}(\phi_{p}(\mathcal{P}^{u}),\phi_{p}(\mathcal{P}^{a}))\\ \end{split}$
@@
$\mathcal{D}_{ua}=\{(\texttt{$x^{u}$}_{j},\texttt{p}^{a}_{j})\}_{j=1}^{N}$, $\mathcal{D}_{uu}=\{(\texttt{$x^{u}$}_{j},\texttt{p}^{u}_{j})\}_{j=1}^{N}$, $\mathcal{P}^{u}=\{\texttt{p}^{u}_{j}\}$, and $\mathcal{P}^{a}=\{\texttt{p}^{a}_{j}\}$, and $\alpha_{\cdot}$ are regularization hyperparameters. We assign equal weight to each of the loss terms, thus choosing $\alpha_{(.)}=1$. The above objective function encourages the CLIP embeddings of $x^{u}$ and $\texttt{p}^{u}$, and $x^{a}$ and $\texttt{p}^{a}$, to be closer together while increasing the distance between $x^{u}$ and $\texttt{p}^{a}$, and $x^{a}$ and $\texttt{p}^{u}$, respectively. During fine-tuning of Espresso, there is a trade-off between effectiveness and utility which is inherent to all other fine-tuning CRTs as well. We subject our fine-tuning to the constraint that achieved the highest effectiveness for the least drop in utility.
We use prompts for fine-tuning in Equation (2) and Equation (3) instead of only concepts. This is because $\texttt{p}^{u}$ and $\texttt{p}^{a}$, containing $c^{u}$ and $c^{a}$, provide more context.
The prompts used for the T2I model are not known before-hand, but we can define the concepts to be filtered. Since the prompt includes these concepts, filtering the images generated from the prompts is sufficient.
Replacement of Blocked Images. As with other filters, on detecting an image with an unacceptable concept, Espresso generates a replacement image as in SDv1.4 (Radford et al., 2021).
5. Experimental Setup
We describe the baselines for CRTs and attacks (Section 5.1), different concepts used for evaluation (Section 5.2), dataset and model configuration (Section 5.3), and metrics (Section 5.4).
5.1. Evaluation Baselines
We compare Espresso with prior state-of-the-art CRTs across R1, R2, and R3. We use six fine-tuning CRTs (CA, FMN, SA, ESD, UCE, and SDD) and one filtering CRT (UD) which are described in Section 2.2.
To evaluate R3, we consider different state-of-art attacks from literature (Section 2.3). We modify (indicated with “+”) existing naïve attacks, Ring-a-Bell, and PEZ, to account for CRTs. The CRTs use some variant of the following optimization: detach $c^{u}$ from p such that the generated image is far from $c^{u}$ and closer to some $c^{a}$. Hence, to design an effective attack which accounts for such CRTs, in addition to the attacks’ original objectives, we minimize the loss between $\texttt{p}^{adv}$ and $c^{a}$ while increasing its loss with $c^{u}$. This, in turn, will move $\phi_{p}(\texttt{p}^{adv})$ closer to $\phi_{p}(c^{u})$ and further from $\phi_{p}(c^{a})$. We modify the attacks by considering the following loss function:
(4) $\begin{split}\mathcal{L}_{att+}=\mathcal{L}_{att}-\alpha_{u}\mathcal{L}_{MSE}(\phi_{p}(c^{u}),\phi_{p}(\texttt{p}^{adv}))+\\ \alpha_{a}\mathcal{L}_{MSE}(\phi_{p}(c^{a}),\phi_{p}(\texttt{p}^{adv}))\end{split}$
@@ 5.1. Evaluation Baselines
The typographic attack (Noever and Noever, 2021) against CLIP is where text characters are superimposed onto an (unrelated) image to fool CLIP by forcing it to focus on the this text instead of the image. We turn this into an attack against CRTs by superimposing $c^{a}$ at the bottom of $x^{u}$. Using the resulting adversarial images, we use PEZ+ to find their corresponding $\texttt{p}^{adv}{s}$. We call this attack Typo+.
5.2. Concepts
Concept Types. We use the same set of unacceptable concepts as in prior work (Kumari et al., 2023a; Pham et al., 2024a; Kim et al., 2023), categorizing them into three groups:
  • Group-1 covers inappropriate concepts such as nudity (e.g., female breasts or male/female genitalia), violence (e.g., bloody scenes, fighting, burning, hanging, weapons, and wars), disturbing (e.g., distorted faces, bodies, human flesh, or bodily fluids), and hateful (e.g., defamatory depictions of races, harmful stereotypes, or Holocaust scenes).

### v4

@@ 3 Problem Statement
Adversary Model. We assume (blackbox) access to a deployed target T2I model (f) to which a client can send an input (p) to generate an image ($x$). Further, f uses some CRT. The goal of the adversary ($\mathcal{A}dv$) is to force f to generate $x^{u}$ despite the presence of a CRT to suppress it. We give an advantage to $\mathcal{A}dv$ by allowing whitebox access to a local identical copy of f with the CRT for use in designing attacks. This is reasonable as $\epsilon_{\theta}$ and CLIP are publicly available. For filtering CRTs, we assume that $\mathcal{A}dv$ has whitebox access to the filter to use its loss function in designing the attacks.
@@
Requirements An ideal CRT should be: R1 Effectivein minimizing the generation of $x^{u}$; R2 Utility-preserving, maintaining the quality of acceptable images (for fine-tuning CRTs) or not blocking them (for filtering CRTs); and R3 Robustto resist evasion with adversarial prompts ($\texttt{p}^{adv}$).
@@
Limitations of Prior Works. In Section 6, we empirically show the limitations of prior works by evaluating R1, R2, and R3 of different CRTs. We provide a brief summary here. Fine-tuning CRTs modify f thereby explicitly creating a trade-off between effectiveness (R1) and utility (R2[16, 30, 68, 21, 21, 19, 26]. Further, these CRTs do not consider robustness (R3) in their design and are susceptible to evasion by $\mathcal{A}dv$ with $\texttt{p}^{adv}$. Filtering CRTs [46, 67] detect unacceptable concepts either in p (aka prompt filter) or in $x$ and block them (a.k.a image filter). Since, they do not modify f, they can maintain utility without impacting effectiveness. Prior filtering approaches may not be accurate in detecting unacceptable concepts (impacts R1[34]. They can also be easily evaded (harms R3) ([48] and Section 6.2). Further, the current state-of-the-art filter, UD [67], trains specialized classifiers for each concept on substantial data, which limits their generalization to new concepts.
4 Espresso: Robust Filtering CRT
@@
We present Espresso, a robust concept filtering CRT which aims to satisfy the requirements from Section 3 and achieve a better trade-off than prior CRTs. Espresso uses a classifier F to detect images and filter unacceptable concepts in generated images. Following SDv1.4, on detecting an unacceptable concept, Espresso generates a replacement image [46, 48]. We identify CLIP as the natural choice for such a classifier as it is 1. pre-trained on a large dataset covering a wide range of concepts, and 2. used across many T2I models, and encodes similar information as seen in them. Hence, CLIP is a better choice for a filter than training specialized classifiers for each concept (e.g., [67]).
@@
However, simply using CLIP for Espresso is not sufficient as seen in Stable Diffusion’s filter ($\texttt{F}_{SD}$[46]. $\texttt{F}_{SD}$ thresholds the cosine similarity between the embeddings of $x$ and each pre-defined unacceptable concept to identify $x^{u}$. Rando et al. [48] design adversarial prompts ($\texttt{p}^{adv}$) to evade $\texttt{F}_{SD}$. Since $\texttt{F}_{SD}$ uses only the cosine similarity to $c^{u}$, it gives $\mathcal{A}dv$ the freedom to modify the prompt embeddings in any direction, while constructing $\texttt{p}^{adv}$, to force a misclassification. We address this in Espresso by (a) configuring CLIP’s classification objective for effectiveness and robustness, and (b) fine-tuning to further improve utility.
@@
Configuring CLIP’s Classification Objective. Instead of using the cosine similarity to only $c^{u}$ as in $\texttt{F}_{SD}$, we configure the objective function of Espresso for filtering $x^{u}$ by using the cosine similarity to both $c^{u}$ and $c^{a}$. Further, jointly optimizing for two embeddings yields better utility as observed in prior fine-tuning CRT[30, 56, 15].
@@
$=\text{argmax}_{i,i\in{a,u}}\left\{\frac{\text{exp}{(\text{cos}(\phi_{x}(x),\phi_{p}(c_{i}))/\tau)}}{\sum_{j\in\{a,u\}}\text{exp}(\text{cos}(\phi_{x}(x),\phi_{p}(c_{j}))/\tau)}\right\}$ (1)
where $\tau=\frac{1}{100}$ is the default temperature parameter used in CLIP. Further, the cosine similarity is equivalent to the scaled Euclidean distance between the embeddings [35]:
@@
$\implies\text{cos}(\texttt{$x$},c)=-\frac{1}{2}(\overline{\phi_{x}(\texttt{$x$})}-\overline{\phi_{p}(c)})^{T}(\overline{\phi_{x}(\texttt{$x$})}-\overline{\phi_{p}(c)})+1$
Hence, by using CLIP as a filter, we project the image embeddings onto the vector between $c^{u}$ and $c^{a}$ and classify it based on its distance to both. This simple configuration change in CLIP’s classification objective restricts $\mathcal{A}dv$ to adding noise only along the low-dimensional vector joining $c^{u}$ and $c^{a}$: $\phi_{p}(c^{u})-\phi_{p}(c^{a})$ in the direction of $c^{a}$. This is more challenging for $\mathcal{A}dv$ to evade than $\texttt{F}_{SD}$, as projecting to a lower-dimensional representation can improve robustness [59, 4]. Hence, we conjecture that this can help satisfy R1 and R3 (as shown later in Section 6).
Fine-tuning. The above configuration change may impact the utility on some concepts due to the close proximity of some $c^{u}$ and $c^{a}$. To restore any drop in utility, we use fine-tuning. We use two different fine-tuning variants depending on the group of concepts.
For the case where $c^{u}$ and $c^{a}$ are opposites (e.g., Group-1 concepts where $c^{u}$ = violence and $c^{a}$ = peaceful), the above objective function might have a low correlation between $\phi_{p}(\texttt{p})$ and the corresponding $\phi_{x}(\texttt{$x$})$. This might result in poor utility. We use following objective function:
$\begin{split}\mathcal{L}_{\textsc{Espresso}}=\alpha_{\text{aa}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{aa})-\alpha_{\text{ua}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{ua})\\ +\alpha_{\text{uu}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{uu})-\alpha_{\text{au}}\mathcal{L}_{\text{Con}}(\mathcal{D}_{au})\\ +\alpha_{\text{uu-t}}\mathcal{L}_{MSE}(\phi_{p}(\mathcal{P}^{u}),\phi_{p}(\mathcal{P}^{a}))\\ \end{split}$ (2)
@@
$\mathcal{D}_{ua}=\{(\texttt{$x^{u}$}_{j},\texttt{p}^{a}_{j})\}_{j=1}^{N}$, $\mathcal{D}_{uu}=\{(\texttt{$x^{u}$}_{j},\texttt{p}^{u}_{j})\}_{j=1}^{N}$, $\mathcal{P}^{u}=\{\texttt{p}^{u}_{j}\}$, and $\mathcal{P}^{a}=\{\texttt{p}^{a}_{j}\}$, and $\alpha_{\cdot}$ are regularization hyperparameters. We assign equal weight to each of the loss terms, thus choosing $\alpha_{(.)}=1$. The above objective function encourages the CLIP embeddings of $x^{u}$ and $\texttt{p}^{u}$, and $x^{a}$ and $\texttt{p}^{a}$, to be closer together while increasing the distance between $x^{u}$ and $\texttt{p}^{a}$, and $x^{a}$ and $\texttt{p}^{u}$, respectively.
In contrast, for concepts in Group-2 and 3, we fine-tune F to maximize the distance between $\phi_{p}(\texttt{p}^{u})$ and $\phi_{p}(\texttt{p}^{a})$, and hence, minimize the following loss:
$L_{\textsc{Espresso}}=-||\overline{\phi_{p}(\texttt{p}^{u})-\phi_{p}(\texttt{p}^{a})}||_{2}$ (3)
We use prompts ($\texttt{p}^{u}$ and $\texttt{p}^{a}$) for fine-tuning in Equation 3 and Equation 2 instead of only concepts ($c^{u}$ and $c^{a}$): $\texttt{p}^{u}$ and $\texttt{p}^{a}$ already contain $c^{u}$ and $c^{a}$, and provide more context.
During fine-tuning of Espresso, there is a trade-off between effectiveness and utility which is inherent to all other fine-tuning CRTs as well. We subject our fine-tuning to the constraint that achieved the highest effectiveness for the least drop in utility. We empirically show the benefit of fine-tuning in Appendix .5: Table XII).
5 Experimental Setup
We use Stable Diffusion v1.4 [49] and its default configuration as the target T2I model following prior CRT[16, 30, 68, 21, 19, 26]. We describe the attack baselines against CRTs (Section 5.1), metrics (Section 5.2), different concepts used for evaluation (Section 5.3), and pipeline used for evaluating CRTs (Section 5.4).
5.1 Revisiting Attack Baselines
@@
$\begin{split}\mathcal{L}_{att+}=\mathcal{L}_{att}-\alpha_{u}\mathcal{L}_{MSE}(\phi_{p}(c^{u}),\phi_{p}(\texttt{p}^{adv}))+\\ \alpha_{a}\mathcal{L}_{MSE}(\phi_{p}(c^{a}),\phi_{p}(\texttt{p}^{adv}))\end{split}$ (4)
[…]
Finally, typographic attack [41] against CLIP is where text characters are superimposed onto an (unrelated) image to fool CLIP by forcing it to focus on this text instead of the image. We turn this into an attack against CRTs by superimposing $c^{a}$ at the bottom of $x^{u}$. Using the resulting adversarial images, we use PEZ+ to find their corresponding $\texttt{p}^{adv}$. We call this attack Typo+.
5.2 Metrics
We now describe the metrics to evaluate each of the requirements. We assume access to a reference CLIP, separate from f with the CRTs. The reference CLIP is a Stable Diffusion v1.4 model, a standard assumption in prior work [30, 16, 19, 26].
R1 (Effectiveness). Depending on the CRT, we use:
  • CLIP accuracy [22, 30] for fine-tuning CRTs. This is the cosine similarity (divided by temperature parameter $\tau^{\prime}=\frac{1}{100}$ of the reference CLIP) between the embeddings of the generated image $x$ with the embeddings of $c^{u}$ and $c^{a}$ from the reference CLIP. This outputs the likelihood of predicting $c^{u}$. Hence, CLIP accuracy should be low (ideally zero) for effective concept removal. Formally, it is:
    $\frac{\text{exp}{(\text{cos}(\tilde{\phi_{x}}(\texttt{$x$}),\tilde{\phi_{p}}(c^{u})/\tau^{\prime}))}}{\text{exp}(\text{cos}(\tilde{\phi_{x}}(\texttt{$x$}),\tilde{\phi_{p}}(c^{u})/\tau^{\prime}))+\text{exp}(\text{cos}(\tilde{\phi_{x}}(\texttt{$x$}),\tilde{\phi_{p}}(c^{a})/\tau^{\prime}))}$
    where $\tilde{\phi_{p}}$ and $\tilde{\phi_{x}}$ are embeddings from the reference CLIP. For filtering CRT, if $x^{u}$ is detected, a replacement image is generated. Here, we calculate the CLIP accuracy on the final set of images after filtering.
  • False Negative Rates (FNR) [67, 48] for filtering CRTs is the fraction of images with $c^{u}$ which are not blocked. It should be low (ideally zero).
R2 (Utility). Depending on the CRT, we use:
  • Normalized CLIP score for fine-tuning CRTs is is the ratio of cosine similarity between $\phi_{x}(\texttt{$x$})$ and $\phi_{p}(\texttt{p})$ from f, compared to that from a reference CLIP as a baseline which is assumed to have the maximum achievable CLIP score. Formally, $\frac{\text{cos}(\phi_{x}(\texttt{$x$}),\phi_{p}(\texttt{p}))}{\text{cos}(\tilde{\phi}_{x}({\texttt{$x$}}),\tilde{\phi}_{p}(\texttt{p}))}$. Normalized CLIP score should be high (ideally one) for high utility. For the normalized CLIP score to be greater than one, the fine-tuned T2I model should generate images that match their prompts more closely than the original T2I model. This is unlikely to occur in practice as seen in our evaluation (c.f. Table VI). This metric is different from standard CLIP score from prior work [22, 30, 26, 19] which measures the cosine similarity between $\phi_{x}(\texttt{$x$})$ and $\phi_{p}(\texttt{p})$ from f. We did this to compare the utility of the T2I with the CRT to that of the T2I without the CRT. Hence, a perfect match between them will give a normalized CLIP score of one.
  • False Positive Rates (FPR) [67, 48] for filtering CRTs is the fraction of images without $c^{u}$ which are blocked. It should be low (ideally zero).
R3 (Robustness). We use the same metrics as effectiveness: CLIP accuracy to evaluate fine-tuning CRTs and compare Espresso with them, and FNR for filtering CRTs.
5.3 Concept Types
We use the same set of unacceptable concepts as in prior work [30, 42, 26], categorizing them into three groups:
  • Group-1 covers inappropriate concepts such as nudity (e.g., female breasts or male/female genitalia), violence (e.g., bloody scenes, fighting, burning, hanging, weapons, and wars), disturbing (e.g., distorted faces, bodies, human flesh, or bodily fluids), and hateful (e.g., defamatory racial depiction, harmful stereotypes, or Holocaust scenes).
@@ 5.3 Concept Types
5.4 Pipeline for Evaluating CRTs
We now describe the pipeline which includes identifying acceptable concepts (step 1), generation of datasets (step 2), training filters or fine-tuning T2I models with CRTs (step 3), and validating acceptable concepts (step 4), and evaluating different CRTs (step 5). We present an overview of the pipeline in Figure 1 and describe each of the steps below.
Figure 1: Step 1 $\rightarrow$ Identify some possible $c^{a}$ for each $c^{u}$. Step 2 $\rightarrow$ Generate datasets containing prompts and corresponding images for both $c^{a}$ and $c^{u}$ for train ($\mathcal{D}_{tr}$), test ($\mathcal{D}_{te}$), and validation ($\mathcal{D}_{val}$). Step 3 $\rightarrow$ Use $\mathcal{D}_{tr}$ to fine-tune T2I model with CRTs and train the filtering classifier. Step 4 $\rightarrow$ Evaluate effectiveness and utility on $\mathcal{D}_{val}$ and choose the best candidate for $c^{a}$. Step 5 $\rightarrow$ Evaluate using $\mathcal{D}_{te}$: for effectiveness (R1), use CLIP accuracy or FNR on unacceptable prompts ($\mathcal{D}_{te}^{u}$), utility (R2) on acceptable prompts ($\mathcal{D}_{te}^{a}$), and robustness (R3) using CLIP score or FPR on $\mathcal{D}_{adv}^{u}$ which includes $\texttt{p}^{adv}$ constructed using attacks on $\mathcal{D}_{te}^{u}$. [Color Scheme and Component Descriptions] $\rightarrow$ Prompts, images and arrows corresponding to unacceptable (acceptable) are indicated in red (green). Metrics are indicated in dashed boxes.
Step 1: Identifying Acceptable Concepts
A good choice of $c^{a}$ is one which effectively steers a T2I model away from generating $x^{u}$ while maintaining utility on other concepts. Hence, the choice of $c^{a}$ can impact the overall effectiveness and utility of CRTs. We select $c^{a}$ for a given $c^{u}$ such that it is either opposite to $c^{u}$ (Group-1) or is a semantic generalization of $c^{u}$ so as to avoid infringing copyrights (Group-2,3). For $c^{u}$ in Group-1, we consider multiple alternative synonyms for $c^{a}$ from which we choose the best possible candidate by measuring effectiveness and utility on a validation dataset (c.f. Step 4). For $c^{u}$ in Group-2,3, we use $c^{a}$ chosen by prior works [30, 21] for a fair comparison. We indicate them in the format “$c^{u}$rightarrow $c^{a}$”:
[…]
  • Group-2: $c^{a}$ is the type of $c^{u}$, following prior works [30, 21, 16]. For instance, Grumpy Cat $\rightarrow$ cat, Nemo $\rightarrow$ fish, Captain Marvel $\rightarrow$ female superhero, Snoopy $\rightarrow$ dog, and R2D2 $\rightarrow$ robot.
[…]
Step 2: Generate and Split Datasets
We describe the generation of train, validation, and test datasets.
Train Datasets. We summarize the different training/fine-tuning dataset configurations across CRTs in Table I.
TABLE I: Training dataset configurations required by CRTs.
CRT Configuration Requirements
CA [30] 200 unacceptable prompts ($\texttt{p}^{a}$) from ChatGPT with one image/prompt from SDv1.4
ESD [16] Unacceptable concept ($c^{u}$)
FMN [68] 8 unacceptabe prompts ($\texttt{p}^{u}$) and one image/prompt from SDv1.4
SDD [26] Unacceptable concept ($c^{u}$) and 10 corresponding images
SA [21] Acceptable concept ($c^{a}$) and 1000 corresponding images, and 6 unacceptable prompts ($\texttt{p}^{u}$)
UCE [19] Unacceptable and acceptable concepts ($c^{u}$ and $c^{a}$)
UD [67] 776 total images with 580 acceptable images ($x^{a}$), and 196 unacceptable images ($x^{u}$): nudity (48), violence (45), disturbing (68), and hateful (35)
Fine-tuning CRTs. Different fine-tuning CRTs use different dataset configurations to ensure that R1 and R2 are met. We use the exact same configuration as described in each of the prior works to ensure that they satisfy the requirements claimed in their papers. While the configuration across different CRTs vary, the configuration for a specific CRT across different concepts is the same, in accordance with the instructions in the respective works. We briefly summarize them in Table I. CA uses acceptable prompts which are generated from ChatGPT such that they contain $c^{a}$, and 1 image per prompt. ESD and SDD both use an unacceptable concept, and SDD additionally uses 10 images corresponding to $c^{u}$. All images are generated using SD v1.4. FMN uses unacceptable prompts of the form “An image of {$c^{u}$}”, and 1 image per prompt. For all the fine-tuning CRTs, we use their publicly available code for training to ensure the configuration is same as reported in their papers [31, 17, 69, 18, 20, 27, 45].
Filtering CRTs. We train $\texttt{F}_{\texttt{UD}}$ using its exact dataset configuration (Table I) and code [67]. For Espresso, we follow the template by Kumari et al. [30] to generate 10 unacceptable and acceptable prompts using ChatGPT, with one image per prompt from SDv1.4. For Group-2 concepts, we randomly select 10 ChatGPT-generated prompts from the datasets provided by Kumari et al. [30]. Our choice for having a small amount of training data is inspired from prior works on poisoning attacks which show that little data is required to modify CLIP [64, 7]. Our results also show that using such little data is indeed sufficient to meet R1, R2, and R3 (c.f. Section 6).
Validation Datasets. We consider the following validation datasets ($\mathcal{D}_{val}$): we denote dataset with acceptable prompts ($\texttt{p}^{a}$) and corresponding images ($x^{a}$) as $\mathcal{D}_{val}^{u}$; and dataset with acceptable prompts ($\texttt{p}^{a}$) and corresponding images ($x^{a}$) as $\mathcal{D}_{val}^{a}$. For R1, we use $\mathcal{D}_{val}^{u}$ by generating 10 unacceptable prompts using ChatGPT, and generate 5 images per prompt using SD v1.4. For R2, we use $\mathcal{D}_{val}^{a}$ by randomly choosing 100 non-overlapping acceptable prompts from the COCO 2014 dataset [36], and 1 image per prompt, generated from SD v1.4. We summarize them in Table II.
TABLE II: Validation dataset configuration.
Concepts Data Configuration
Group-1 (R1) $\mathcal{D}_{val}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 5 images ($x^{u}$) per prompt from SDv1.4
Group-2 (R1) $\mathcal{D}_{val}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 5 images ($x^{u}$) per prompt from SDv1.4
Group-3 (R1) $\mathcal{D}_{val}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 5 images ($x^{u}$) per prompt from SDv1.4
All (R2) $\mathcal{D}_{val}^{a}$ 100 acceptable prompts ($\texttt{p}^{a}$) and 1 image ($x^{a}$) per prompt from the COCO 2014 dataset
Test Datasets. We consider test datasets ($\mathcal{D}_{te}$): we denote the dataset with acceptable prompts ($\texttt{p}^{a}$) and corresponding images ($x^{a}$) as $\mathcal{D}_{te}^{a}$; and with unacceptable prompts ($\texttt{p}^{u}$) and corresponding images ($x^{u}$) as $\mathcal{D}_{te}^{u}$.
TABLE III: Evaluation dataset Configuration.
Concepts Data Configuration
Group-1 (R1, R3) $\mathcal{D}_{te}^{u}$ I2P Dataset [52] of unacceptable prompts: for nudity: 449 prompts, violence: 758, disturbing: 857, hateful: 235
Group-2 (R1, R3) $\mathcal{D}_{te}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 20 images ($x^{u}$) per prompt from SDv1.4
Group-3 (R1, R3) $\mathcal{D}_{te}^{u}$ 10 unacceptable prompts ($\texttt{p}^{u}$) from ChatGPT and 20 images ($x^{u}$) per prompt from SDv1.4
All (R2) $\mathcal{D}_{te}^{a}$ 200 acceptable prompts ($\texttt{p}^{a}$) from COCO 2014 dataset
For evaluating effectiveness (R1) and robustness (R3), we use $\mathcal{D}_{te}^{u}$ which is generated as follows: For Group-1, we use an independent benchmark dataset, Inappropriate Image Prompts (I2P), containing prompts likely to generate unsafe images [52]. It covers all four Group-1 concepts. For nudity, I2P includes a “nudity percentage” attribute. We choose all unacceptable prompts with a nudity percentage $>10$, resulting in a total of 300 unacceptable prompts, in accordance with prior works [16, 19]. UD [67] also used political as a concept, which is excluded from our evaluation since it is not a part of I2P. Note that I2P prompts do not explicitly contain $c^{u}$, i.e., the words nude, violent, hateful, or disturbing are not explicitly included in the prompts. For Group-2 and Group-3 concepts, there are no standard benchmark datasets. Hence, we generate the dataset by following prior works [30, 16] to obtain 200 unacceptable images generated from SDv1.4 from 10 unacceptable prompts generated from ChatGPT. For R2, we use $\mathcal{D}_{te}^{a}$ which includes the COCO 2014 test dataset with 200 randomly chosen acceptable prompts. We summarize the datasets in Table III.
Step 3: Fine-tune T2I model with CRTs or train filter
Using $\mathcal{D}_{tr}$, we fine-tune T2I models using different fine-tuning CRTs whose optimizations are described in Section 2.3. We use six fine-tuning CRTs (CA, FMN, SA, ESD, UCE, and SDD). For filtering CRT, we consider one baseline (UD) and do not compare with SD-Filter as UD outperforms it. For UD, we train the filtering classifier on their dataset. We fine-tune Espresso on $\mathcal{D}_{tr}$ and use the CLIP L-patch-14 [1] due to its popularity, which is also the default text encoder with Stable Diffusion v1.4. However, other variants of CLIP are also applicable.
Step 4: Select best $c^{a}$ using $\mathcal{D}_{val}$